OpenStudy (anonymous):

The last Question of the day A rectangle is inscribed in the semi-circle of radius r units with side lying in the diameter of the semi - circle. The maximum possible area of the rectangle in terms of radius r is....

OpenStudy (anonymous):

 r^2=x^2+h^2 AREA of the rectangle=2x*h